# Distributionally Robust Learning for Multi-source Unsupervised Domain Adaptation

### v4

@@ 3.3 Covariate Shift Setting
For each source $l\in[L]$, suppose that ${\mathbf{Q}}_{X}$ is absolutely continuous with respect to $\mathbf{P}^{(l)}_{X}$ and define the density ratio $\omega^{(l)}(x)=d{\mathbf{Q}}_{X}(x)/d\mathbf{P}^{(l)}_{X}(x)$. A crucial step in our approach is to estimate these density ratios $\{\omega^{(l)}\}_{l\in[L]}$. To ensure independence between the density ratio estimators and the data used for bias correction, we employ the sample-splitting technique, analogous to that used for constructing the individual source model estimators $\{\widehat{f}^{(l)}\}_{l\in[L]}$ in (3.5). Specifically, we construct the density ratio estimator $\widehat{\omega}^{(l)}_{\mathcal{A}}$ using the covariate data from the split $\mathcal{A}_{l}$ combined with the target covariate $\{X^{\mathbf{Q}}_{j}\}_{j\in[N]}$, and $\widehat{\omega}^{(l)}_{\mathcal{B}}$ using the data from $\mathcal{B}_{l}$ together with the target covariate data. In Supplements’ Section A.4 [53], we provide details on the construction of $\widehat{\omega}^{(l)}_{\mathcal{A}}$ and $\widehat{\omega}^{(l)}_{\mathcal{B}}$ via the Bayes formula. Our proposal is also compatible with alternative density ratio estimation methods [22, 18, 35, e.g.,]. We then define the sample-split density ratio estimator $\widehat{\omega}^{(l)}(X^{(l)}_{i})$ for a sample $X^{(l)}_{i}$ from the $l$-th source domain:
@@ 4.1 Rate Improvement with Bias Correction
The first part of Assumption 3 imposes a boundedness condition on $w^{(l)}$, which is standard in the density ratio estimation literature (see, e.g., Assumptions 13.1 and 14.7 in [49]). This condition is also consistent with the positivity (or overlap) assumption commonly invoked in causal inference [39, 10, e.g.]. The second part captures the convergence rate of the density ratio estimators $\{\widehat{\omega}^{(l)}\}_{l\in[L]}$ via the sequence $\eta_{\omega}$, where $\eta_{\omega}$ may or may not converge to 00.

### v5

@@ 3.3 Covariate Shift Setting
For each source $l\in[L]$, suppose that ${\mathbf{Q}}_{X}$ is absolutely continuous with respect to $\mathbf{P}^{(l)}_{X}$ and define the density ratio $\omega^{(l)}(x)=d{\mathbf{Q}}_{X}(x)/d\mathbf{P}^{(l)}_{X}(x)$. A crucial step in our approach is to estimate these density ratios $\{\omega^{(l)}\}_{l\in[L]}$. To ensure independence between the density ratio estimators and the data used for bias correction, we employ the sample-splitting technique, analogous to that used for constructing the individual source model estimators $\{\widehat{f}^{(l)}\}_{l\in[L]}$ in (3.5). Specifically, we construct the density ratio estimator $\widehat{\omega}^{(l)}_{\mathcal{A}}$ using the covariate data from the split $\mathcal{A}_{l}$ combined with the target covariate $\{X^{\mathbf{Q}}_{j}\}_{j\in[N]}$, and $\widehat{\omega}^{(l)}_{\mathcal{B}}$ using the data from $\mathcal{B}_{l}$ together with the target covariate data. In Supplements’ Section A.4 [62], we provide details on the construction of $\widehat{\omega}^{(l)}_{\mathcal{A}}$ and $\widehat{\omega}^{(l)}_{\mathcal{B}}$ via the Bayes formula. Our proposal is also compatible with alternative density ratio estimation methods [25, 20, 40, e.g.,]. We then define the sample-split density ratio estimator $\widehat{\omega}^{(l)}(X^{(l)}_{i})$ for a sample $X^{(l)}_{i}$ from the $l$-th source domain:
@@ 4.1 Rate Improvement with Bias Correction
The first part of Assumption 3 imposes a boundedness condition on $w^{(l)}$, which is standard in the density ratio estimation literature (see, e.g., Assumptions 13.1 and 14.7 in [56]). This condition is also consistent with the positivity (or overlap) assumption commonly invoked in causal inference [45, 12, e.g.]. The second part captures the convergence rate of the density ratio estimators $\{\widehat{\omega}^{(l)}\}_{l\in[L]}$ via the sequence $\eta_{\omega}$, where $\eta_{\omega}$ may or may not converge to 00.